### v2

# Computing hydration free energies of small molecules with first principles accuracy

Abstract
Free energies play a central role in characterising the behaviour of chemical systems and are among the most important quantities that can be calculated by molecular dynamics simulations. The free energy of hydration in particular is a well-studied physicochemical property of drug-like molecules and is commonly used to assess and optimise the accuracy of nonbonded parameters in empirical forcefields, and also as a fast-to-compute surrogate of performance for protein-ligand binding free energy estimation. Machine learned potentials (MLPs) show great promise as more accurate alternatives to empirical forcefields, but are not readily decomposed into physically motivated functional forms, which has thus far rendered them incompatible with standard alchemical free energy methods that manipulate individual pairwise interaction terms. However, since the accuracy of free energy calculations is highly sensitive to the forcefield, this is a key area in which MLPs have the potential to address the shortcomings of empirical forcefields. In this work, we introduce an efficient alchemical free energy method compatible with MLPs, enabling, for the first time, calculations of biomolecular free energy differences with ab initio accuracy. Using a pretrained, transferrable, alchemically equipped MACE model, we demonstrate sub-chemical accuracy for the hydration free energies of organic molecules.
preprint: AIP/123-QEDfootnotetext: jhm72@cam.ac.uk
I Introduction
Free energies are arguably the most important quantities accessible by molecular simulation in soft matter and play an important role in computational drug discovery [1, 2]. Free energy methods rely on rigorous thermodynamic protocols, enabling prediction of protein-ligand binding [3, 4], small molecule solvation and solubility [5], crystal polymorphs of organic molecules[6] and protein residue mutation [7].
Driven by the widespread availability of efficient GPU hardware, highly optimised molecular dynamics (MD) software [8] and carefully parameterised forcefields [9, 10], free energy calculations have become an industry standard tool, accounting for a large fraction of the computation performed in pharmaceutical R&D. Relative binding free energy (RBFE) calculations in particular have emerged as a crucial aspect of structure-based drug discovery, helping to screen and rank congeneric series of compounds during hit-to-lead and lead optimisation campaigns [1].
Whilst various methods for calculating free energies for small molecules via MD have been proposed, alchemical transformations have emerged as the de facto standard approach, allowing free energy estimates to be obtained in 1-2 GPU hours per compound [11, 12] Rather than explicitly simulating the a compound undergoing a change of state as it would happen in reality (which would be computationally prohibitive), alchemical calculations take advantage of the fact that the free energy is a state function and so the free energy difference between two state points is independent of the pathway connecting them. By introducing an alchemical parameter, $\lambda$, a new Hamiltonian is constructed as a linear combination of the Hamiltonians describing the two end states,
$H(\vec{r},\lambda)=\lambda H_{1}(\vec{r},\lambda)+(1-\lambda)H_{0}(\vec{r},\lambda)$ (1)
The free energy difference associated with the transformation can then be computed, for example by thermodynamic integration
[…]
where $\langle\dots\rangle_{\lambda}$ denotes the ensemble average corresponding to the Hamiltonian $H(\lambda)$. Using such artificial pathways leads to enormous computational savings, since the sampling required to explore intermediate state points is orders of magnitude less than it would be for a real pathway describing the physical process. In practice, the transformation is accomplished by identifying the atoms that change during the transformation (i.e. those atoms whose chemical element changes, or that disappear entirely) and interpolating the corresponding forcefield parameters between the end states. This requires the use of so-called dummy atoms to model atoms that disappear during the transformation. These atoms maintain their bonded terms but have their interactions with the surroundings switched off. This allows for transformations like morphing a hydrogen atom to a methyl group or isolating a molecule from its environment, while minimally perturbing the phase space required to be explored by the system.
Early attempts to perform free energy calculations with MD faced energy conservation and convergence issues due to the divergence of the nonbonded energy (usually a Lennard-Jones potential) when partially decoupled atoms overlap [13, 14]. Beutler et al were the first to address this issue by incorporating softening parameters to scale nonbonded interactions as a function of the alchemical parameter $\lambda$, thus averting singularities in the potential energy as atoms come into contact [15]
$U(\lambda,r)=4\epsilon\lambda^{n}\Bigg{[}\Bigg{.}\left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-2}-\\ \left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-1}\Bigg{.}\Bigg{]}$ (3)
where $\epsilon$ and $\sigma$ are the Lennard-Jones well depth and radius, respectively, and $\alpha_{LJ}$ and $m$ are positive, tunable constants that control the smoothness of the softcore function. Although Beutler-type soft core potentials are commonly used in various simulation packages, it is worth noting that several alternative approaches have been suggested to further improve numerical stability and address specific failure modes [16, 17], and alternative functional forms with a more natural soft core have also been tried [18].
While soft-core potentials have been widely successful, a significant limitation of current state-of-the-art free energy calculations lies in the accuracy of the empirical forcefields. Specifically, the constraints imposed by the functional form, omission of energetic contributions and restricted atom-type parameters in small molecule forcefields have a significant impact on accuracy [1]. Particularly notable are the widespread use of fixed-charge electrostatic models that do not account for geometry-dependent polarization, and the limited accuracy of torsional barriers, which often require refitting to bespoke DFT calculations of each new compound to achieve reasonable accuracy [19, 20, 21].
In recent years, transferable machine learned potentials (MLP)s have been introduced as a compelling alternative to empirical forcefields, demonstrating significant advances in materials modelling and biomolecular simulation [22, 23, 24, 25, 26, 27]. Emerging from Behler and Parinello’s initial work [28], the field has expanded rapidly, with many architectures being proposed to accurately model the QM potential energy surface. A particularly successful innovation has been the specific encoding of permutational, rotatational and translational symmetries [29]. These approaches have led to a series of data-efficient, universal potentials for biomolecular and materials simulation, enabling accurate prediction of atomistic and thermodynamic properties for a wide range of chemical compositions, albeit at an increased cost compared to empirical potentials.
Although MLPs can deliver significant accuracy improvements compared to empirical forcefields, their computational expense and the high sampling requirements of free energy calculations have meant that their application has been mostly restricted to corrective perturbations [30, 31, 32]. In this approach, the intramolecular interactions of a subset of atoms (usually the organic molecule) are modelled by the MLP, while the remaining bonded interactions and all nonbonded interactions are modelled by the empirical forcefield. Although there is some evidence that this approach can modestly improve the accuracy of RBFE calculations, it does not address the parametrisation of nonbonded interactions, which are known to be crucial for obtaining accurate free energies [31]. Indeed, it has recently been shown that, for hydration free energies, MLP corrections with mechanical embedding fail to provide statistically significant accuracy improvements [32]. It is worth noting that recently more sophisticated electrostatic embedding ML/MM schemes have also been proposed. However, the performance of these methods in corrective binding free energy calculations has not yet been established [33].
In parallel, a handful of free energy methods that do not rely on alchemical transformations have been proposed, making the calculation conceptually more straightforward, easier to set up and more amenable to MLPs. One promising example uses the alchemical transfer method (ATM) [34], which relies on a coordinate transformation to interpolate between two physical end states [35, 36]. It has been shown that, especially in combination with the ANI-2x potential for modelling ligand intramolecular forces [22], this method is competitive with the commercially available FEP+ package [3], which uses the empirical OPLS4 forcefield [37]. This work still relies upon a mechanically embedded ML/MM Hamiltonian, and since both physical end states must be accommodated within the same simulation box, it is less efficient than alchemical perturbation methods due to the large amount of excess solvent required [34].
Given the potential for significant accuracy improvements, there is a need for rigorous and efficient alchemical free energy methods that are compatible with (i) describing the entire system using MLPs, and (ii) existing well-tested free energy protocols and analysis packages (e.g. pymbar [38], pmx [39, 40]) that are currently in use with classical forcefields. In this work, we present a pre-trained, transferable biomolecular MLP based on the recently introduced MACE-OFF23 potentials [23] that is uniquely equipped with scalable soft-core interactions, enabling numerically stable simulations of condensed phase systems with alchemically decoupled intermolecular interactions. We provide an efficient implementation of the method in the widely used OpenMM package [8]. Furthermore, we demonstrate the fast convergence and sub-chemical accuracy of our approach in the calculation of hydration free energies for a series of small organic molecules. To the best of our knowledge, this work represents the first thermodynamically rigorous, ab initio-accurate approach to condensed phase hydration free energy calculations.
II Results & Discussion
II.1 Alchemical simulations with machine learned forcefields
To address the divergence of the potential energy predicted by a MLP fitted to DFT as atoms overlap when sampling intermediate values of $\lambda$, we augment the training set with dimer configurations containing artificially softened two-body interactions to avoid the formal energy divergence. We further incorporate a $\lambda$ dependence via modification of the non-trainable parameters of the model architecture to scale nonbonded interactions between the solute and the solvent. This combined approach results in $\lambda$-dependent many-body interactions that enable smooth and regularised interpolation between the coupled and decoupled end states.
II.1.1 Softcore dimer curves
To construct the softcore dimer curves, we started by matching the value and gradient of the DFT force with a quartic function of the form $ax^{4}+b$ that approaches a constant value as separation goes to zero (Figure 1). The softcore energies were then obtained by integration to ensure data consistency. The transition point was adjusted for each dimer curve to minimise the gradient at the switching point, which in turn limited the maximum energy of pairwise interactions to a numerically stable value (see Figure 1). Importantly, the transition point was chosen so as not to alter the position and curvature of the energy minimum, preserving the equilibrium properties of the potential. This also ensured that the softcore region is only explored during alchemical simulations in which the nonbonded interactions are (partially) decoupled, since it lies several eV higher in energy than the equilibrium.
The sharper curvature of quartic functions enabled smaller maximum forces and energies compared to lower-order polynomials that have been employed in previous softcore formulations [16]. Using the original MACE-OFF23 dataset augmented with the synthetic dimer curve, we then fitted the softcore potential from scratch. Since it is expected that the learned radial embeddings will differ substantially from the MACE-OFF23 models due to the new softcore data, we found that this resulted in lower force and energy errors compared to transfer learning from the original MACE-OFF23 checkpoints.
Figure 1: Construction of softcore dimer curve for the Br–N diatomic pair by matching the gradient at a set switching point (in this example, 1.73 Å) to that given by DFT.
The errors of the test set on the original MACE-OFF23 data set were comparable to the performance of the original potentials, indicating that the model’s ability to learn energies and forces of equilibrium configurations is not affected by additional soft-core data (section V.1).
II.1.2 Alchemical modification to MACE
Whilst in principle the softcore formulation described above is sufficient to perform stable alchemical free energy calculations, in practice, the resulting linear formulation, where the functional form of the potential has no explicit dependence on $\lambda$, can lead to significant challenges with phase space overlap [41]. Functions lacking this $\lambda$ dependence typically require complex reweighting strategies to adjust window spacing on the fly to maintain reasonable overlap between neighbouring replicas. To avoid these issues, we also modified the nonlearnable parameters in MACE to enable alchemical scaling of selected nonbonded interactions, resulting in a formulation analogous to classical soft core force fields  [15, 16, 17]. By modifying only the nonlearnable components of MACE, our approach enables alchemical simulations of arbitrary systems without additional fine tuning of the model.
Whilst we introduce alchemical modifications in terms of the MACE architecture, we stress that our approach can, in principle, be applied to any architecture that utilises the ‘density trick’ to construct atomic environment vectors without significant modification [42].
Here we describe the changes to the MACE architecture required to implement scaled nonbonded interactions. For a full description of the architecture, we direct the reader to the original publications [43, 29].
MACE predicts the total energy of a system as a sum of atom-wise contributions. Each atom’s environment is constructed from the relative coordinates and atomic numbers of its neighbours, up to a fixed cutoff. These high body-order atomic features are efficiently constructed by taking tensor products of two-body terms that describe pairwise interactions.
For each atom, the intial features of its neighbours are combined with the interatomic displacement vectors, $\bm{r}_{ij}$, to form the one-particle basis $\phi_{ij,k\eta_{1}l_{3}m_{3}}^{(t)}$. The radial distance is used as an input into a learnable radial function $R(r_{ij})$ with several outputs that correspond to the different ways in which the displacement vector and the node features can be combined while preserving equivariance [44]:
$\phi_{ij,k\eta_{1}l_{3}m_{3}}^{(t)}(\lambda)$ $=\sum_{l_{1}l_{2}m_{1}m_{2}}C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}$ (4)
$\alpha_{ij}R_{k\eta_{1}l_{1}l_{2}l_{3}}^{(t)}(r_{ij})\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$
where $Y_{l}^{m}$ are the spherical harmonics, and $C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}$ denotes the Clebsch-Gordan coefficients. There are multiple ways of constructing an equivariant combination with a given symmetry $(l_{3},m_{3})$, and these multiplicities are enumerated by the path index $\eta_{1}$ [45, 43].
In a similar approach to other recently published work [46], our formulation modifies MACE by including an additional factor of $\alpha_{ij}$ that scales selected two-body terms in the one-particle basis. These terms are selected as those that cross the alchemical boundary, for example edges connecting atoms in the solute and solvent in the case of a hydration free energy calculation.
$\alpha_{ij}=\begin{cases}\lambda&\text{if }i\in\text{solute}\oplus j\in\text{solute}\\ 1&\text{else}\end{cases}$ (5)
The one-particle basis functions $\phi_{ij,k\eta_{1}l_{3}m_{3}}^{(t)}$ are then passed through multiple layers of message passing, during which tensor products are taken to construct the many-body symmetric features.
${\bm{B}}^{(t),\nu}_{i,\eta_{\nu}kLM}=\sum_{{\bm{l}}{\bm{m}}}\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}\prod_{\xi=1}^{\nu}A_{i,kl_{\xi}m_{\xi}}^{(t)}$ (6)
This approach (named MACE-OFF23-SC) enables smoothly scaled many-body nonbonded interactions whilst maintaining the correct physical description of the end states, where the decoupled system’s graph is identical to that of the separated components, ensuring consistent free energy calculations.
In MACE-OFF23-SC, an approximately linear $\lambda$ schedule can be applied to the alchemical transformation (Figure 2), regardless of the size of the perturbation, compared to the highly skewed schedule required for linear softcore potentials [41]. It is also worth noting that, unlike linear scaling schemes, this approach is amenable to nonequilibrium switching approaches, in which $\lambda$ is driven at a constant rate between the end states [4, 47].
While this approach can be applied in principle to any many-body potential that constructs site-wise energies as a tensor product of two-body features, we note that the success of this approach depends on the strong regularisation of MACE to enable smooth interpolation between the end states, as the contributions to the node energies are a nonlinear function of the (scaled) two-body features.
Figure 2: Dependence on $\lambda$ of softcore two-body interactions for a C–F diatomic pair learned by MACE-OFF23-SC.
We note that, whilst our alchemical modifications to the MACE graph share similarities with recently published work by Nam et al [46], the additional inclusion of the softcore potential is crucial for stable alchemical condensed phase simulations, where, unlike materials applications, direct atomic overlap in partially decoupled states must be handled.
II.2 Hydration free energy calculations
Figure 3: Comparison of MACE-OFF23-SC hydration free energies with classical forcefields and experiment
Hydration free energies play a crucial role in assessing the accuracy of the intermolecular interactions within a forcefield and are an essential component of the standard evaluation of thermodynamic property prediction. These simulations probe the subtle nonbonded interactions between molecules, but their sampling requirements remain small relative to, for example, protein–ligand binding free energies, hence errors can be primarily attributed to the accuracy of the forcefield.
However, unlike properties such as densities and enthalpies of vapourisation and mixing, which are directly observable as ensemble averages from MD trajectories, free energies require evaluation of expectation values such as that shown in Eq. (2). Here we employ an alternative formulation, based on Bennett’s Acceptance Ratio (BAR) [48, 38], for the free energy difference between states A and B, given by
$\Delta G_{AB}=-k_{b}T\,\ln{\left\langle e^{-\beta(U_{A}(\vec{r})-U_{B}(\vec{r}))}\right\rangle_{A}}$ (7)
[…]
These calculations can be efficiently performed using perturbations in parameter space, where the intermolecular forces are gradually decoupled between the solute and solvent. Although such free energy calculations are routine with empirical forcefields, to the best of our knowledge, this work demonstrates the first rigorous, ab initio quality hydration free energy calculations.
Figure 4: Convergence of ethane hydration free energy with simulation time.
II.2.1 Organic molecule benchmark
We first evaluated the ability of MACE-OFF23-SC to compute accurate, converged hydration free energies of a series of small organic molecules (Figure 3). For this test, we selected compounds from the FreeSolv database that cover both polar and nonpolar molecules, in addition to several examples known to be challenging for small molecule forcefields [49]. We performed replica exchange molecular dynamics with an ensemble of replicas with $\lambda$ windows spaced in the interval [0,1] to connect the physical end states. To expedite phase space exploration, exchanges of replicas between adjacent $\lambda$ windows were attempted every 1 ps.
For the small systems, we found that 12 replicas enabled sufficient phase-space overlap to converge the free energy simulations (Figure 5). For caffeine and ibuprofen, 20 replicas were required for sufficient exchange probabilities. The replica spacing was hand-tuned from an initial linear spacing to concentrate sampling in the region $0.15\leq\lambda\leq 0.4$, where the curvature of $\frac{\partial H}{\partial\lambda}$ is greatest. For each system, convergence of the free energy as a function of time was confirmed (Figure 4).
Figure 5: Transition probability matrix between 12 replicas from REMD simulation of methanol in water.
Across the range of molecules studied here, we find that free energy predictions computed using MACE-OFF23-SC match or, in many cases, outperform widely-used classical forcefields, with approximately a factor of 2 improvement in MAE and RMSE and improved correlation with experiment (Table 1).
MAE RMSE $r^{2}$
MACE-OFF23-SC 0.53 0.60 0.99
GAFF 2.04 2.59 0.74
OpenFF 2.1 1.79 2.24 0.81
Table 1: Summary of hydration free energy errors (kcal/mol).
For example, compared to OpenFF and GAFF, we find that the MACE-OFF23-SC prediction for methanol falls within the experimental error, whilst the classical forcefields underpredict the magnitude of the hydration free energy. This trend has previously been studied in the context of empirical forcefields and has been attributed to the assignment of partial charges on aliphatic hydroxyl groups [50]. Since our approach is not dependent upon the coarse graining of chemical space through atom-typing, our approach is not susceptible to the same inaccuracies.
To further evaluate the performance of our method, we computed the hydration free energy of several larger molecules, including ibuprofen and caffeine. While still small and rigid compared to a typical therapeutic small-molecule, they are closer to drug-like chemical space than the small organic molecules studied thus far. Due to the larger excluded volume, 20 replicas were required for sufficient phase space overlap, however our method converged quickly to within experimental error. This is particularly encouraging given that, for caffeine, both OpenFF and GAFF predictions have MAEs of around 5 kcal/mol, suggesting that this approach might scale to larger, conformationally flexible drug-like molecules at a similar level of accuracy.
III Conclusion
We have introduced a modified version of the MACE-OFF23 general organic machine learned forcefield that enables, for the first time, theoretically rigorous condensed phase alchemical free energy simulations with first principles based machine learned forcefields. A dual approach of softcore repulsion and $\lambda$-dependent two-body interactions enable numerically well-behaved simulations of weakly interacting atoms, directly analogous to softcore formulations of classical forcefields. Our method enables access to quantities directly relevant to drug discovery at ab initio quality.
We tested the utility of our approach by calculating hydration free energies for a series of small organic molecules. Our method achieves sub-chemical accuracy for a range of polar and nonpolar compounds, improving on state-of-the-art empirical forcefield errors by a factor of two. Equally importantly, we demonstrate that convergence of the free energy is fully feasible within reasonable simulation time, even for the challenging case of disappearing an entire molecule.
We envisage soft-core-equipped machine-learning potentials becoming an increasingly exploited avenue for accurate and efficient calculation of free energies. As well as being an important quantity in forcefield benchmarking, free energies represent key optimisation quantities in drug discovery, where they can give direct access to protein–ligand binding affinities. Hence ab initio quality computational methods are an important tool when rigorous and accurate predictions are required. We further expect that these methods will form part of the benchmarking process for new machine learned forcefields, particularly those incorporating explicit treatment of long range interactions.
IV Methods
IV.1 MACE-OFF23-SC fitting
In previous work, it was found that the use of equivariant features ($L_{max}>0$) in the tensor product significantly increased the accuracy of intermolecular force predictions compared to the invariant MACE-OFF23 model. All simulations in this work were performed using a model with $L_{max}=1$ and 64 channels, which we found to adequately balance computational performance and accuracy.
IV.2 Equilibrium free energy perturbation
Equilibrium free energy calculations were performed using OpenMM, and made use of modified versions of the openmmtools, openmm-ml and openmm-torch libraries to perform Hamiltonian replica exchange and interface MACE with OpenMM [8, 51, 52, 53, 54, 55]. This approach enabled interpolation between the coupled and decoupled endstates via a $\lambda$ dependent Hamiltonian.https://github.com/jharrymoore/mace-md
Since MACE predicts total intermolecular interactions without distinguishing between repulsive exchange and attractive Coulomb and dispersion terms, solvent-solute interactions are switched off using a single transformation, parameterised by the alchemical parameter $\lambda$. We adopt a direct decoupling approach in which the organic molecule is decoupled from a box of water, whilst all other interactions are unmodified. This results in a decoupled end state equivalent to a separated water box and small molecule in vacuum.
Initial solvated structures were generated with pdbfixer, using a rhombic dodecahedral box and a padding distance of 2.0 nm, as required by the 10 Å receptive field of MACE. Structures were energy minimised using the L-BFGS algorithm, and subsequently equilibrated in the NPT ensemble for 50 ps. The last frame was used to seed the replica exchange calculations.
The phase space perturbation was spanned by 12 replicas, with the following spacing [0.0, 0.12, 0.16, 0.20, 0.2425, 0.28, 0.4, 0.52, 0.64, 0.76, 0.88, 1.0], resulting in at least tridiagonal overlap in all cases. Hamiltonian replica exchange between all replicas were attempted every 1000 steps of dynamics (1 ps). It was found that 1 ns of sampling was sufficient to converge free energy estimates in all cases.
Replica exchange calculations were performed using ReplicaExchangeSampler functionality from openmmtools. Each simulation was performed on a single node, comprising 4 NVIDIA A100 GPUs. Replicas were assigned a single MPI rank, and 3 MPI ranks were mapped to each GPU.
IV.3 Empirical forcefield solvation free energies
GAFF solvation free energies and standard errors were taken from the FreeSolv Dataset [49]. OpenFF values were calculated using OpenFF 2.1.0 via the Absolv code, using the default nonequilbrium switching protocolhttps://github.com/SimonBoothroyd/absolv/tree/main.
Acknowledgements
JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb for their helpful discussions. DJC acknowledges support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). We acknowledge computational resources provided by the Cambridge Service for Data-Driven Discovery (CSD3), the AstraZeneca Scientific Computing Platform (SCP) and the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1
Competing interests
GC has equity stake in Symmetric Group LLP and Ångstrom AI Inc., companies engaged in the application of machine learning to material and molecular simulation.